# LLM-Meta-SR: Learning to Evolve Selection Operators for Symbolic Regression

### v1

@@ Abstract
Large language models (LLMs) have revolutionized algorithm development, yet their application in symbolic regression, where algorithms automatically discover symbolic expressions from data, remains constrained and is typically designed manually by human experts. In this paper, we propose a learning-to-evolve framework that enables LLMs to automatically design selection operators for evolutionary symbolic regression algorithms. We first identify two key limitations in existing LLM-based algorithm evolution techniques: code bloat and a lack of semantic guidance. Bloat results in unnecessarily complex components, and the absence of semantic awareness can lead to ineffective exchange of useful code components, both of which can reduce the interpretability of the designed algorithm or hinder evolutionary learning progress. To address these issues, we enhance the LLM-based evolution framework for meta symbolic regression with two key innovations: bloat control and a complementary, semantics-aware selection operator. Additionally, we embed domain knowledge into the prompt, enabling the LLM to generate more effective and contextually relevant selection operators. Our experimental results on symbolic regression benchmarks show that LLMs can devise selection operators that outperform nine expert-designed baselines, achieving state-of-the-art performance. This demonstrates that LLMs can exceed expert-level algorithm design for symbolic regression.
@@ 1 Introduction
Symbolic regression is the task of discovering mathematical expressions that accurately model a given dataset. It is widely used in domains such as finance [54], education [53],chemistry [9], and physics [15], due to its interpretability and ability to uncover underlying relationships. Formally, given a dataset $\mathcal{D}=\{\mathbf{x}^{(i)},y^{(i)}\}_{i=1}^{N}$, where $\mathbf{x}^{(i)}\in\mathbb{R}^{d}$ represents input features and $y^{(i)}\in\mathbb{R}$ denotes the corresponding target value, the goal is to find a symbolic expression $f^{*}\in\mathcal{F}$ from a space of candidate expressions $\mathcal{F}$ that minimizes a predefined loss function $\mathcal{L}$, i.e.,
@@
Numerous efforts have been devoted to designing symbolic regression algorithms. These approaches can be broadly categorized as evolutionary symbolic regression [5, 17], neural symbolic regression [3, 27], and search-based symbolic regression [65, 70] approaches. In recent years, there has been growing interest in hybrid methods that combine the flexibility of evolutionary algorithms with the learning capabilities of neural networks [32, 22, 71].
In both evolutionary [5, 17] and neural-evolutionary [32, 22, 71] symbolic regression algorithms, the system follows an iterative optimization process in which solutions are progressively refined by selecting promising candidates. In this process, the selection operator responsible for identifying candidate solutions worth refining, plays a key role in system effectiveness. However, in current evolutionary and neural-evolutionary symbolic regression systems, the selection operator is typically manually designed by domain experts. This manual process requires significant effort and trial-and-error, which can limit the pace of progress. Ideally, an automated algorithm design framework capable of generating effective selection operators for symbolic regression would substantially enhance research and development efficiency in this area.
Large language models (LLMs) have recently emerged as powerful tools for automated heuristic design in various optimization tasks [52, 35, 69]. However, most existing frameworks primarily focus on generating heuristic rules for specific combinatorial optimization tasks. In this paper, we pursue a more ambitious goal: to automatically design a core algorithmic component—specifically, the selection operator—in evolutionary symbolic regression, with the aim of achieving strong performance across a wide range of regression tasks. While frameworks such as FunSearch [52], EoH [35], and ReEvo [69] provide useful foundations for algorithm evolution, two key challenges remain underexplored in LLM-driven algorithm evolution. The first is complexity bias, illustrated in Figure 2. LLMs tend to generate overly long or intricate code during the optimization process, reflecting a phenomenon analogous to code bloat in genetic programming [2]. This results in the accumulation of redundant or non-functional logic, which impairs interpretability, wastes a large number of tokens, and slows optimization. As shown in Figure 2, such complexity can lead to performance stagnation, ultimately limiting the effectiveness of evolution.
Figure 1: Average code length of solutions in the population over generations, along with the score of the best solution.
@@
The second is the underutilization of semantic information in the generated code. Here, we define semantic information as the performance of an algorithm on individual task instances. Typically, only the average performance across all training instances is provided to the LLM to guide code generation [23]. However, this average does not sufficiently capture the fine-grained behavior of the algorithm across different tasks. As shown in Figure 2, consider a scenario with three algorithms: the first two perform well on the first dataset but poorly on the second, while the third shows the opposite pattern. Despite their contrasting behaviors, all three algorithms may exhibit similar average performance. However, presenting the first two algorithms together to the LLM offers limited benefit, as they exhibit redundant patterns—both succeeding and failing on the same datasets. In contrast, combining the first and third algorithms allows the LLM to integrate complementary strengths, potentially generating an algorithm that performs well on both datasets. Thus, incorporating semantic information is crucial for effective algorithm evolution.
@@
  • We propose an LLM-driven meta symbolic regression framework to automatically design selection operators. This learning-to-evolve process enables the discovery of generalizable operators that consistently outperform human-designed counterparts on a wide range of unseen datasets.
@@ 2.1 Selection Operators in Symbolic Regression
There has been significant research on designing selection operators for symbolic regression, including tournament selection [63], Boltzmann sampling [56, 52], lexicase selection [30], novelty search [33], and complementary phenotype selection [14]. These operators aim to improve symbolic regression algorithms from different perspectives. Lexicase selection highlights the importance of specialists on subsets of training instances [30]. Novelty search emphasizes diversity by explicitly encouraging solutions that behave differently from one another to escape local optima [33]. Complementary phenotype selection focuses on leveraging complementary individuals in crossover and using this information during the selection process [14]. Despite these advances, there remains a lack of selection operators that simultaneously capture multiple desirable traits. A key reason is that designing such operators is labor-intensive and relies heavily on trial and error.
@@ 2.2 LLMs for Symbolic Regression
Early work on LLMs for symbolic regression focused on specialized language models [27, 3], often assisted by Monte Carlo Tree Search [28, 55] or evolutionary algorithms [71]. With the advent of general-purpose LLMs, their use in symbolic regression has attracted increasing attention [56], either through the FunSearch framework [52] or via integration with evolutionary algorithm frameworks [22]. In these studies, LLMs effectively replace traditional crossover and mutation operators to generate candidate solutions, but selection operators are still manually designed. This motivates the exploration of automated selection operator design using LLMs.
Some recent work incorporates dataset descriptions into symbolic regression, referring to this as “semantics” [37]. However, this differs from the concept of semantics in genetic programming [43], where it refers to the behavior of a program. In LLM-based program evolution, the fine-grained behavior of candidate programs is often overlooked and reduced to aggregate scores, which can obscure meaningful information and hinder evolutionary progress.
@@ 3 Algorithm Framework
This work introduces a learning-to-evolve framework that leverages large language models (LLMs) within a meta symbolic regression setting to automate the design of selection operators. The key components are as follows.
@@ 3.1 Solution Representation
Each solution is a piece of code representing a selection operator. The input to the selection operator is a list of individuals. Each individual contains a list representing squared error, a list of predicted values, and a list of nodes, which can be used to compute the height and depth of the symbolic tree. The evolutionary status, specifically the ratio between the current generation number and the total number of generations, is also provided. The expected output of the selection operator is a list of promising symbolic trees. The structure of the selection operator is presented in LABEL:lst:template of the supplementary material.
@@ 3.2 Meta-Evolution Workflow
The meta-evolution algorithm consists of two nested loops: an outer meta-evolution loop and an inner symbolic regression loop. The outer loop is responsible for generating new selection operators, while the inner loop evaluates their performance. An overview of the workflow is shown in Figure 3.
@@
  • Population Initialization: The initial population $\mathcal{P}^{(0)}$ is generated by prompting the LLM to produce $N$ random selection operators. The details of the initialization prompts are provided in Appendix A of the supplementary material.
  • Synthetic Evaluation: Since LLM-generated code may contain syntax errors or infinite loops, we run code on several basic test cases to filter out invalid operators. Each $O_{i}^{(t)}\in\mathcal{P}^{(t)}$ is evaluated in a synthetic testing environment and discarded if it contains syntax/runtime errors, exceeds a 300-second execution time, or is more than 100$\times$ slower than the fastest operator in the batch. The test cases are provided in Appendix C of the supplementary material.
  • Solution Evaluation: For each selection operator $O_{i}^{(t)}\in\mathcal{P}^{(t)}$, its fitness is computed as $f(O_{i}^{(t)})=\frac{1}{T}\sum_{j=1}^{T}\texttt{SR}(O_{i}^{(t)},\mathcal{D}_{j}),$ where $\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ are symbolic regression datasets, and $\texttt{SR}(O_{i}^{(t)},\mathcal{D}_{j})$ denotes the symbolic regression performance of $O_{i}^{(t)}$ on dataset $\mathcal{D}_{j}$. In brief, each dataset is split into a training and validation set. Symbolic regression is performed on the training set, and the fitness of the selection operator is computed based on its performance on the validation set. Details of the solution evaluation procedure are provided in Appendix D of the supplementary material.
@@
  • Solution Generation: Based on the selected parents, $N$ candidate operators are generated per generation. Of these, $N-M$ are generated via crossover and $M$ via mutation:
    • Operator Crossover: For each pair $(O_{a}^{(t)},O_{b}^{(t)})$, a new operator is generated as $O_{\text{new}}^{(t+1)}=\texttt{LLM\_Crossover}(O_{a}^{(t)},O_{b}^{(t)})$. The goal is to generate a new piece of code that combines effective building blocks from the selected parent operators, guided by their performance scores on the evaluated datasets.
    • Operator Mutation: Let $O^{*}=\arg\max_{O_{i}^{(t)}\in\mathcal{P}^{(t)}}f(O_{i}^{(t)})$ be the best-performing operator in generation $t$ [69]. Mutated variants are generated as $O_{\text{mut}}^{(t+1)}=\texttt{LLM\_Mutate}(O^{*})$, with the objective of generating novel code based on the elite operator.
    The prompts used for crossover and mutation are provided in Appendix A.
@@
3.3 Bloat Control
Prompt-based Length Limit: To mitigate code bloat in the evolution of selection operators, we incorporate a length constraint in the prompt during solution generation. For any operator $O_{i}\in\mathcal{P}^{(t)}$, its code length is denoted by $\ell(O_{i})$, measured as the number of non-empty, non-comment lines in the implementation. Rather than enforcing a hard upper bound $\ell(O_{i})\leq L_{\max}$, which is difficult to control during LLM generation, we embed the maximum length $\ell_{\text{target}}$ directly into the prompt. For instance, the LLM is instructed: “Write a selection operator with code length $\leq\ell_{\text{target}}$.” This prompt-guided strategy encourages the model to produce more concise programs.
Multi-Objective Survival Selection: Beyond prompt-based constraints, we employ a multi-objective survival selection strategy based on both operator fitness and code length to further control bloat, since LLMs do not always follow instructions to generate code within a specified length. Each operator $O_{i}$ is represented as a tuple $(f(O_{i}),\ell(O_{i}))$, where $f(O_{i})$ denotes its average task performance across $T$ tasks, and $\ell(O_{i})$ denotes its code length. The parents for generating the next generation $\mathcal{P}^{(t+1)}$ are selected from the combined set $\mathcal{P}^{(t)}\cup\mathcal{P}^{\text{offspring}}$ using a dominance-dissimilarity selection mechanism [67].
The key idea is to compute a dominance score for each operator based on weak Pareto dominance [34] and code similarity. An operator $O_{i}$ is said to weakly Pareto dominate another operator $O_{j}$, denoted $O_{i}\succeq O_{j}$, if and only if $f(O_{i})\geq f(O_{j})$ and $\ell(O_{i})\leq\ell(O_{j})$. For each such pair $(O_{i},O_{j})$, the score of $O_{j}$ is penalized by the code similarity $\text{sim}(O_{i},O_{j})$, computed using the CodeBLEU metric [51]. The total dominance score of $O_{j}$ is then defined as $s(O_{j})=\sum_{O_{i}\succeq O_{j}}-\text{sim}(O_{i},O_{j})$. Operators are then ranked, and the top-$N$ operators with higher scores are selected, as they are less frequently dominated and exhibit greater dissimilarity from dominating counterparts. This selection strategy encourages a trade-off between maximizing task performance and minimizing code complexity, effectively reducing operator bloat while maintaining diversity and effectiveness.
@@ 3.4 Semantics-Aware Evolution
To this end, we propose a semantics-aware selection strategy, with the pseudocode provided in Algorithm 1. Because each selection operator $O_{i}\in\mathcal{P}^{(t)}$ has been evaluated across $d$ datasets and is associated with a score vector $\mathbf{s}_{i}=[s_{i,1},\dots,s_{i,d}]$, we can compute a complementarity score for each candidate in the population. The process begins by randomly selecting the first parent $O_{a}\in\mathcal{P}^{(t)}$. Random selection, as used in ReEvo [69] and HSEvo [12], helps mitigate premature convergence [12]. Then, for each candidate $O_{i}\in\mathcal{P}^{(t)}$, we compute a complementarity score:
@@
which measures the potential combined performance of $O_{a}$ and $O_{i}$ across all datasets. The second parent $O_{b}$ is selected as the operator with the highest complementarity score, i.e., $O_{b}=O_{i^{*}}$ where $i^{*}=\arg\max_{i}\mu_{i}$.
@@
0:  $\mathcal{P}^{(t)}=\{O_{1},\dots,O_{N}\}$: population where each $O_{i}$ has score vector $\mathbf{s}_{i}$
1:  $O_{a}\leftarrow$ Random sample from $\mathcal{P}^{(t)}$

### v2

# LLM-Meta-SR: In-Context Learning for Evolving Selection Operators in Symbolic Regression

@@ Abstract
Large language models (LLMs) have revolutionized algorithm development, yet their application in symbolic regression, where algorithms automatically discover symbolic expressions from data, remains constrained and is typically designed manually by human experts. In this paper, we propose a meta learning framework that enables LLMs to automatically design selection operators for evolutionary symbolic regression algorithms. We first identify two key limitations in existing LLM-based algorithm evolution techniques: a lack of semantic guidance and code bloat. The absence of semantic awareness can lead to ineffective exchange of useful code components, and bloat results in unnecessarily complex components, both of which can reduce the interpretability of the designed algorithm or hinder evolutionary learning progress. To address these issues, we enhance the LLM-based evolution framework for meta symbolic regression with two key innovations: a complementary, semantics-aware selection operator and bloat control. Additionally, we embed domain knowledge into the prompt, enabling the LLM to generate more effective and contextually relevant selection operators. Our experimental results on symbolic regression benchmarks show that LLMs can devise selection operators that outperform nine expert-designed baselines, achieving state-of-the-art performance. Moreover, the evolved operator can further improve the state-of-the-art symbolic regression algorithm, achieving the best performance among 26 symbolic regression and machine learning algorithms across 116 regression datasets. This demonstrates that LLMs can exceed expert-level algorithm design for symbolic regression.
@@ 1 Introduction
Symbolic regression (SR) is the task of discovering mathematical expressions that accurately model a given dataset. It is widely used in domains such as finance (Shi et al. 2025), education (Shen et al. 2024),chemistry (Chen et al. 2025), and physics (Feng et al. 2025), due to its interpretability and ability to uncover underlying relationships. Formally, given a dataset $\mathcal{D}=\{\mathbf{x}^{(i)},y^{(i)}\}_{i=1}^{N}$, where $\mathbf{x}^{(i)}\in\mathbb{R}^{d}$ are input features and $y^{(i)}\in\mathbb{R}$ are target values, the goal is to find a symbolic expression $f^{*}\in\mathcal{F}$ from a space of candidate expressions $\mathcal{F}$ that minimizes a loss function $\mathcal{L}$:
@@
Many SR algorithms have been proposed, including evolutionary (Jiang and Xue 2024; Fong and Motani 2024b), neural (Biggio et al. 2021; Kamienny et al. 2022), and search-based (Kahlmeyer, Fischer, and Giesen 2025; Yu et al. 2025) approaches. In recent years, there has been growing interest in hybrid methods that combine the flexibility of evolutionary algorithms with the learning capabilities of neural networks (Landajuela et al. 2022; Grayeli et al. 2024; Zhang et al. 2025).
In both evolutionary (Jiang and Xue 2024; Fong and Motani 2024b) and neural-evolutionary (Landajuela et al. 2022; Grayeli et al. 2024; Zhang et al. 2025) symbolic regression algorithms, the system follows an iterative optimization process in which solutions are progressively refined by selecting promising candidates. In this process, the selection operator responsible for identifying candidate solutions worth refining, plays a key role in system effectiveness. However, current selection operators are typically manually designed by experts, requiring significant effort and trial-and-error, which can limit the pace of progress. Ideally, an automated algorithm design framework capable of generating effective selection operators for symbolic regression would substantially enhance research and development efficiency in this area.
Large language models (LLMs) have recently emerged as powerful tools for automated heuristic design in various optimization tasks (Romera-Paredes et al. 2024; Liu et al. 2024a; Ye et al. 2024). However, most existing frameworks primarily focus on generating heuristic rules for specific combinatorial optimization tasks. In this paper, we pursue a more ambitious goal: to automatically design a core algorithmic component—specifically, the selection operator—in evolutionary symbolic regression, with the aim of achieving strong performance across a wide range of regression tasks. While frameworks such as FunSearch (Romera-Paredes et al. 2024), EoH (Liu et al. 2024a), and ReEvo (Ye et al. 2024) lay groundwork for algorithm evolution, two challenges remain underexplored in LLM-driven algorithm evolution. The first is the underutilization of semantic information in the generated code. Here, we define semantic information as the performance of an algorithm on individual task instances. Typically, only the average performance across all training instances is provided to the LLM to guide code generation (Huang et al. 2025a), which overlooks fine-grained behavior. As illustrated in Figure 2, consider a scenario with three algorithms: the first two perform well on the first dataset but poorly on the second, while the third shows the opposite pattern. Despite their contrasting behaviors, all three algorithms may exhibit similar average performance. However, presenting the first two algorithms together to the LLM offers limited benefit, as they exhibit redundant patterns—both succeeding and failing on the same datasets. In contrast, combining the first and third algorithms allows the LLM to integrate complementary strengths, potentially generating an algorithm that performs well on both datasets. Thus, incorporating semantic information is crucial for effective algorithm evolution.
@@
Figure 2: Average code length of solutions in the population over generations without bloat control, and the score of the best solution.
The second challenge is complexity bias, illustrated in Figure 2. LLMs tend to generate overly long or intricate code during optimization, similar to code bloat in genetic programming (Banzhaf et al. 1998). This results in the accumulation of redundant or non-functional logic, which impairs interpretability, wastes a large number of tokens, and slows optimization. As shown in Figure 2, such complexity can lead to performance stagnation, ultimately limiting the effectiveness of evolution.
@@
  • We propose an LLM-driven meta symbolic regression framework to automatically design selection operators using in-context learning. The LLM learns from design and evaluation history to automatically discover generalizable operators that consistently outperform human-designed counterparts across a wide range of unseen datasets.
@@ 2.1 Selection Operators in Symbolic Regression
There has been significant research on designing selection operators for symbolic regression, including tournament selection (Xie and Zhang 2012), Boltzmann sampling (Shojaee et al. 2025; Romera-Paredes et al. 2024), lexicase selection (La Cava et al. 2019), novelty search (Lehman and Stanley 2010), and complementary phenotype selection (Dolin, Arenas, and Merelo 2002). These operators aim to improve symbolic regression from different perspectives. Lexicase selection highlights the importance of specialists on subsets of training instances (La Cava et al. 2019). Novelty search encourages behavioral diversity to escape local optima (Lehman and Stanley 2010). Complementary phenotype selection focuses on leveraging complementary individuals in crossover and using this information during the selection process (Dolin, Arenas, and Merelo 2002). Despite these advances, there remains a lack of selection operators that simultaneously capture multiple desirable traits, largely because designing such operators remains labor-intensive and trial-and-error-driven.
@@ 2.2 LLMs for Symbolic Regression
Early work on LLMs for symbolic regression focused on specialized language models (Kamienny et al. 2022; Biggio et al. 2021), often assisted by Monte Carlo Tree Search (Kamienny et al. 2023; Shojaee et al. 2024) or evolutionary algorithms (Zhang et al. 2025). With the advent of general-purpose LLMs, their use in symbolic regression has attracted increasing attention (Shojaee et al. 2025), either through the FunSearch framework (Romera-Paredes et al. 2024) or via integration with evolutionary algorithm frameworks (Grayeli et al. 2024). In these studies, LLMs effectively replace traditional crossover and mutation operators to generate candidate solutions, but selection operators are still manually designed. This motivates the exploration of automated selection operator design using LLMs.
Some recent work incorporates dataset descriptions into symbolic regression, referring to this as “semantics” (Liu, Huynh, and van der Schaar 2025). However, this differs from the concept of semantics in genetic programming (Moraglio, Krawiec, and Johnson 2012), where it refers to the behavior of a program. In LLM-based program evolution, the fine-grained behavior of candidate programs is often overlooked and reduced to aggregate scores, which can obscure meaningful information and hinder evolutionary progress.
@@ 3 Algorithm Framework
This work introduces a meta learning framework that leverages large language models (LLMs) within a meta-symbolic regression setting to automate the design of selection operators. The key components are as follows.
@@ 3.1 Solution Representation
Each solution is a piece of code representing a selection operator. The input to the selection operator is a list of individuals. Each individual contains a list representing squared error, a list of predicted values, and a list of nodes, which can be used to compute the height and depth of the symbolic tree. The evolutionary status, specifically the ratio between the current and total generations, is also provided. The expected output of the selection operator is a list of promising symbolic trees. The structure of the selection operator is presented in LABEL:lst:template of the supplementary material.
@@ 3.2 Meta-Evolution Workflow
The meta-evolution algorithm consists of two nested loops: an outer meta-evolution loop and an inner symbolic regression loop. The outer loop generates new selection operators, while the inner loop evaluates their performance. An overview of the workflow is shown in Figure 3.
@@
  • Population Initialization: The initial population $\mathcal{P}^{(0)}$ is generated by prompting the LLM to produce $N$ random selection operators. The details of the initialization prompts are provided in Appendix K of the supplementary material.
[…]
  • Solution Evaluation: For each selection operator $O_{i}^{(t)}\in\mathcal{P}^{(t)}$, its fitness is computed as $f(O_{i}^{(t)})=\frac{1}{T}\sum_{j=1}^{T}\texttt{SR}(O_{i}^{(t)},\mathcal{D}_{j}),$ where $\mathcal{D}_{1},\dots,\mathcal{D}_{T}$ are symbolic regression datasets, and $\texttt{SR}(O_{i}^{(t)},\mathcal{D}_{j})$ denotes the symbolic regression performance of $O_{i}^{(t)}$ on dataset $\mathcal{D}_{j}$. In brief, each dataset is split into a training and validation set. Symbolic regression is performed on the training set, and the fitness of the selection operator is computed based on its performance on the validation set. Details of the solution evaluation procedure are provided in Appendix A of the supplementary material.
@@
  • Solution Generation: Based on the selected parents, $N$ candidate operators are generated per generation. Of these, NMN-M are generated via crossover and $M$ via mutation:
    • Operator Crossover: For each pair $(O_{a}^{(t)},O_{b}^{(t)})$, a new operator is generated as $O_{\text{new}}^{(t+1)}=\operatorname*{LLM\_Crossover}\big{(}O_{a}^{(t)},O_{b}^{(t)}\big{)}$. The goal is to generate a new piece of code that combines effective building blocks from the selected parent operators, guided by their performance scores on the evaluated datasets.
    • Operator Mutation: Let $O^{*}=\arg\max_{O_{i}^{(t)}\in\mathcal{P}^{(t)}}f(O_{i}^{(t)})$ be the best-performing operator in generation $t$ (Ye et al. 2024). Mutated variants are then generated as $O_{\text{mut}}^{(t+1)}=\operatorname*{LLM\_Mutate}(O^{*})$, with the objective of generating novel code based on the elite operator.
    The prompts used for crossover and mutation are provided in Appendix K. Solution generation leverages the in-context learning capability of LLMs (Gao and Das 2024), expecting that LLMs can generate better solutions by analyzing historical execution results.
@@ 3.3 Semantics-Aware Evolution
To this end, we propose a semantics-aware selection strategy, with the pseudocode provided in Algorithm 1. Because each selection operator $O_{i}\in\mathcal{P}^{(t)}$ has been evaluated across $d$ datasets and is associated with a score vector $\mathbf{s}_{i}=[s_{i,1},\dots,s_{i,d}]$, we can compute a complementarity score for each candidate in the population. The process begins by randomly selecting the first parent $O_{a}\in\mathcal{P}^{(t)}$. Random selection, as used in ReEvo (Ye et al. 2024) and HSEvo (Dat, Doan, and Binh 2025), helps mitigate premature convergence (Dat, Doan, and Binh 2025). Then, for each candidate $O_{i}\in\mathcal{P}^{(t)}$, we compute a complementarity score:
@@
which measures the potential combined performance of $O_{a}$ and $O_{i}$ across all datasets. The second parent $O_{b}$ is retrieved as the operator with the highest complementarity score, i.e., $O_{b}=O_{i^{*}}$ where $i^{*}=\arg\max_{i}\mu_{i}$. This process can be viewed as retrieval-augmented generation (RAG), where complementarity serves as the similarity function to retrieve the most complementary code from the population to guide generation.
@@
[…]
@@
3.4 Bloat Control
Prompt-based Length Limit: To mitigate code bloat in the evolution of selection operators, we incorporate a length constraint in the prompt during solution generation. For any operator $O_{i}\in\mathcal{P}^{(t)}$, its code length is denoted by $\ell(O_{i})$, measured as the number of non-empty, non-comment lines in the implementation. We choose to constrain the number of lines, rather than the number of tokens, because line count is less sensitive to variations in variable name length. Rather than enforcing a hard upper bound $\ell(O_{i})\leq L_{\max}$, which is difficult to control during LLM generation, we embed the maximum length $\ell_{\text{target}}$ directly into the prompt. For instance, the LLM is instructed: “Write a selection operator with code length $\leq\ell_{\text{target}}$.” This prompt-guided strategy encourages the model to produce more concise programs.
Multi-Objective Survival Selection: Beyond prompt-based constraints, we employ a multi-objective survival selection strategy based on both operator fitness and code length to further control bloat, since LLMs do not always follow instructions to generate code within a specified length. Each operator $O_{i}$ is represented as a tuple $(f(O_{i}),\ell(O_{i}))$, where $f(O_{i})$ denotes its average task performance across $T$ tasks, and $\ell(O_{i})$ denotes its code length. The parents for generating the next generation $\mathcal{P}^{(t+1)}$ are selected from the combined set $\mathcal{P}^{(t)}\cup\mathcal{P}^{\text{offspring}}$ using a dominance-dissimilarity selection mechanism (Yao et al. 2025).
The key idea is to compute a dominance score for each operator based on weak Pareto dominance (Lin et al. 2022) and code similarity. An operator $O_{i}$ is said to weakly Pareto dominate another operator $O_{j}$, denoted $O_{i}\succeq O_{j}$, if and only if $f(O_{i})\geq f(O_{j})$ and $\ell(O_{i})\leq\ell(O_{j})$. For each such pair $(O_{i},O_{j})$, the score of $O_{j}$ is penalized by the code similarity $\text{sim}(O_{i},O_{j})$, computed using the CodeBLEU metric (Ren et al. 2020). The total dominance score of $O_{j}$ is then defined as $s(O_{j})=\sum_{O_{i}\succeq O_{j}}-\text{sim}(O_{i},O_{j})$. Operators are then ranked, and the top-$N$ operators with higher scores are selected, as they are less frequently dominated and exhibit greater dissimilarity from dominating counterparts. This selection strategy encourages a trade-off between maximizing task performance and minimizing code complexity, effectively reducing operator bloat while maintaining diversity and effectiveness.